d^3y/dx^3-3 d^2y/dx^3+ 4 dy/dx -2y=e^x+cosx


1
Expert's answer
2021-02-24T12:42:04-0500

Given equation is-


"\dfrac{d^3y}{dx^3}-3\dfrac{ d^2y}{dx^3}+ 4 \dfrac{dy}{dx} -2y=e^x+cosx"


It's auxilary equation is:-

"m^3-3m^2+4m-2=0"

"\Rightarrow (m-1)(m^2-2m+2)=0"


The roots of the above equation are "1,1\pm i"


Complimentary function (C.F.) is given by

= "C_1e^x+(C_2cosx+C_3sinx)e^x"

Particular Integral-


"=\dfrac{e^x}{D^3-3D^2+4D-2}+\dfrac{cosx}{D^3-3D^2+4D-2}"


"=\dfrac{xe^x}{3D^2-6D+4}+\dfrac{cosx}{D.(-1^2)-3(-1^2)+4D-2}"


"=\dfrac{xe^x}{3-6+4}+\dfrac{cosx}{3D+1}"


"=\dfrac{xe^x}{1}+(1+3D)^{-1}cosx"


"=xe^x+(1-3D)cosx"


="xe^x+cosx+3sinx"


Therefore Complete solution is-

"y=C.F.+P.I."


"y=C_1e^x+(C_2cosx+C_3sinx)e^x+xe^x+cosx+3sinx"